# Conditional Prototype Rectification Prompt Learning

### v1

@@ 1 Introduction
Leveraging the vast array of image-text association pairs, the trained visual-language models (VLMs) embody critical general knowledge, thereby exhibiting enhanced generalizability across diverse tasks. In recent developments, several visual-language models like Contrastive Language-Image Pre-training [19] (CLIP), Flamingo [1], and ALIGN [12] have been introduced. While VLMs are proficient at extracting visual and textual descriptions, their training necessitates extensive, high-quality datasets. Nevertheless, amassing substantial data for training task-specific models in real-world visual-language applications poses a challenge. To circumvent this issue, efficient transfer learning (ETL) methods [30, 26, 29, 15, 32]have been devised to fine-tune pre-trained VLMs for downstream tasks, demonstrating remarkable efficacy in various few-shot or zero-shot visual recognition tasks.
@@
There are two notable ETL methods for VLMs: prompt tuning [30, 26, 29, 31] and adapter-style tuning [9, 15, 32]. Prompt tuning modifies the textual classifier by incorporating learnable prompts at the input, significantly outperforming zero-shot CLIP with few-shot examples, seen in CoOp [30] and CoCoOp [29]. Conversely, adapter-style tuning adjusts textual classifiers or visual features at the output, using the textual encoder just once. For example, CLIP-Adapter [9] employs a bottleneck layer to fine-tune embeddings, surpassing zero-shot CLIP by 3.02% on ImageNet in a one-shot scenario. TaskRes [27] adjusts textual embeddings with task-specific parameters, acting as independent residuals. Another approach enhances downstream task knowledge using large scale databases or generative models, like Sus-X [23] and C2A [21].
@@
However, two main limitations exist in current ETL research: 1) Many approaches, like CLIP-Adapter [9] and TaskRes [27], focus on task-specific knowledge from only one modality, relying solely on visual or textual features for adaptation. 2) Using generative models or large scale datasets to augment datasets typically requires substantial computational resources, impacting efficiency. In low-data scenarios, limited samples struggle to adequately reveal structural knowledge for downstream tasks, potentially biasing the model toward certain attributes like color or shape. This can limit transferability and generalization. Therefore, integrating multi-modality knowledge and discovering efficient strategies to enrich the dataset are essential for VLM tuning, addressing the biases that stem from data scarcity.
@@
To alleviate these limitations, we propose two prototype rectification strategies, referred to as conditional prototype rectification learning, which aim to model task-specific knowledge for downstream tasks by integrating textual and visual structural knowledge as well as knowledge from unlabeled samples. To tackle the first limitation, we introduce the Conditional Adapter (CoAdapter), which leverages the input image alongside visual and textual prototypes to capture structured knowledge pertinent to downstream tasks, producing sample-specific prototypes. This equips the feature adapter with a comprehensive and accurate understanding essential for these tasks. Addressing the second limitation, we devise the Nearest Neighbor Rectification (NNR) strategy. It identifies the $k$ nearest samples from the unlabeled dataset that align with the prototypes generated by the CoAdapter, facilitating rectification. This process effectively utilizes the latent knowledge within the unlabeled samples, enriching the learning context of model. As shown in Figure 1, CPR achieves significant improvements over the previous state-of-the-art algorithms on both few-shot classification and base-to-new generalization tasks.
@@
  • We introduce a novel adapter-style tuning approach, the Conditional Adapter (CoAdapter), for efficiently transfer learning (ETL) of vision-language models (VLMs). This strategy harnesses dual-modality structure knowledge from both textual and visual domains, allowing the feature adapter to utilize integrated visual and textual insights. This fusion facilitates enhanced learning of task-specific knowledge from downstream tasks, leading to effective VLM tuning.
@@
  • We evaluate our CPR across 11 popular benchmarks in both few-shot classification and base-to-new generalization settings. The evaluations reveal that CPR markedly surpasses earlier prompt-based and adapter-style methodologies, showcasing notable efficacy even in demanding fine-grained image classification tasks, exemplified by FGVCAircraft.
@@
2 Related Works
2.1 Visual-Language Pre-training
@@
The principal aim of Visual-Language Pre-training (VLP) is to execute multimodal pre-training employing extensive datasets comprising image-text pairs, aiming to develop universal representations for text and images. This approach is intended to improve the effectiveness of model in various downstream tasks. These tasks notably include zero-shot learning, few-shot classification, cross-modal generation, open-world segmentation, and open-world detection, among others. In this context, models such as VisualBERT, OSCAR, and Uniter use BERT to process raw text, achieving significant results in multimodal tasks like Visual Question Answering (VQA), which demonstrates their substantial application potential. These successes emphasize the importance of well-designed fusion encoders in merging cross-modal interactions. Additionally, recent studies like CLIP, DeCLIP, and ALIGN have demonstrated the effectiveness of visual-language contrastive learning in producing features that are useful for downstream tasks. These features are acquired by simply calculating the dot product between visual and linguistic embeddings, offering a straightforward understanding of multimodal interactions. Moreover, these approaches avoid the need for self-attention or cross-attention mechanisms, allowing for the precomputation and storage of multimodal embeddings without the need for further modifications, thus increasing efficiency and easing their integration into other tasks.
2.2 Efficient Transfer Learning
@@
The objective of Efficient Transfer Learning (ETL) is to facilitate the transfer of specific task knowledge to downstream tasks by adjusting selected parameters in pre-trained visual-language models (VLMs). ETL can primarily be categorized into two methods: prompt tuning and adapter-style tuning.
Prompt tuning originates from the field of Natural Language Processing (NLP) and is designed to apply differentiated prompt templates for various NLP tasks. To adapt pre-trained VLMs to downstream tasks, the prompt tuning method introduces the use of task-related textual tokens to infer specific textual knowledge for the task. For instance, in the CLIP model, a manually crafted template ”a photo of a [CLASS]” is employed to construct text embeddings for zero-shot prediction, enhancing performance by 1.9% on the ImageNet dataset compared to the linear probing technique. However, these manually defined prompts have limited capability in describing downstream tasks, overlooking the specific knowledge of the current task. To address this issue, Context Optimization (CoOp) introduces learnable soft prompts derived from few-shot annotated examples instead of fixed manual prompts. The limitation of CoOp lies in that it generates a unique and fixed learnable prompt for images in each task, neglecting the variation in image features. Consequently, Conditional Context Optimization (CoCoOp) was developed to generate image-specific contextual prompts, which are combined with textual context for prompt tuning. Specifically, CoCoOp utilizes a lightweight neural network to create a learnable text prompt vector. To acquire tokens highly relevant to the task, ProDA introduces the learning of a prior distribution for prompts. Moreover, ProGrad selectively updates only those prompt gradients that align with the ”general knowledge” generated from the original prompts.
Distinct from prompt tuning, the adapter-style approach modulates textual and visual features to adapt VLMs to downstream tasks. In early work, CLIP-Adapter employed a bottleneck layer to enhance few-shot classification performance. TaskRes [27] introduced a task-independent adapter aimed at decoupling the prior knowledge of pre-trained model from task-specific knowledge. Furthermore, some studies explored acquiring additional prior knowledge from other large pre-trained models, such as DINO, which also showed promising results in the context of adapter-style applications.
@@
In existing methodologies, the works most relevant to our study are CoCoOp and Sus-X. CoCoOp can be regarded as the baseline model for our proposed CPR. Unlike CoCoOp, our CPR considers not only the input image itself but also the structural relationship between the input image and both visual and textual prototypes to further enhance the performance of learnable prompts. Sus-X introduces a text-image generative model to augment the original support set, whereas CPR utilizes unlabeled samples from the test set for prototype refinement, offering greater efficiency.
3 Preliminaries
@@
We first provide brief reviews on the adopted VLM, i.e., contrastive language-image pre-training (CLIP) [19], recap two mainstream approaches for ETL on VLMs, i.e., prompt tuning and adapter-style tuning.
3.1 Contrastive Language-Image Pre-training
@@
Contrastive Language-Image Pre-training (CLIP) [19] aims to acquire visual representations through natural language supervision. CLIP consists of two encoders: one for images and another for text. The image encoder can be either ResNet [10] or Vision Transformer [7] (ViT), transforming images into feature vectors. The text encoder is a Transformer [24] that takes a sequence of word tokens as input and produces a vectorized representation.
@@
CLIP is trained using 400 million image-text pairs and employs a contrastive loss to learn a joint embedding space for both modalities, enabling CLIP to effectively capture a wide range of visual concepts and learn universal visual representations. Specifically, for an input image $x$ belonging to one of the classes $Y=\{y_{1},y_{2},...,y_{C}\}$, the image encoder $\phi$ extracts image features $z=f(x)\in\mathbb{R}^{d}$. To obtain the corresponding text features $w_{i}\in\{1,...,C\}$, names of all given classes are filled into the fixed prompt template ”a photo of a [CLASS]” (e.g., bear, urocissa erythrorhyncha), generating text descriptions $A$, which are then further encoded into text embeddings $W=g(A)\in\mathbb{R}^{d\times C}$. Image-text alignment is optimized based on the cosine similarity of features.
@@
3.2 Revisiting Previous Efficient Transfer Paradigms
@@
CoOp [30] introduces prompt tuning to VLMs for the first time, utilizing $M$ learnable context vectors $v=\{v_{1},v_{2},...,v_{M}\}$to generate task-specific templates, replacing the fixed text prompt templates used in CLIP. Specifically, $t_{i}=\{v_{1},v_{2},...,v_{M},c_{i}\}$ integrates learnable context vectors $v$ with the class token embedding $c_{i}$. This combined prompt is then input into the text encoder $g(\cdot)$. During training process, CoOp maintains the pre-trained VLMs parameters unchanged and solely adjusts the learnable vectors $v$. CoCoOp [29] is an enhanced version of CoOp. CoOp suggests that contexts conditioned on individual instances offer superior generalization because they pivot attention from a narrow set of classes (thus reducing overfitting risks) toward each unique input instance, facilitating broader task applicability. Meanwhile, CoCoOp introduces a Meta-Net to produce a conditional token that is subsequently amalgamated with the context vectors, enhancing adaptability and specificity to each instance.
@@
where $\alpha$ is a scaling factor.In CLIP-Adapter, the adapter module $\phi_{\omega}$ consists of two linear transformation layers with a ReLU activation function situated between them. CLIP-Adapter explores the use of adapters for both visual and textual components, applying the adapter module separately to the image and text branches of CLIP. This investigation reveals that both types of adapters—visual and textual—yield similarly effective performance. When training on downstream tasks, adapter-style approaches like CLIP-Adapter focus on tuning the parameters of these adapter modules only, leaving the foundational pre-trained CLIP model parameters unchanged. This strategy allows for task-specific customization while maintaining the general capabilities of the underlying model. Moreover, recent developments in SuS-X [23] and C2A [21] have utilized generative diffusion models or large-scale external databases to achieve enhanced feature representations $f^{\prime}$.
4 Approach
@@
4.1 Conditional Adapter
Previous works like CoCoOp [29] considered how input images impact text prototypes. And CLIP-Adapter [9], TaskRes [27], and Tip-Adapter [28] derive features using separate visual or textual features. In contrast, we propose conditional adapter (CoAdapter), which leverages the connections between input images and both visual and textual prototypes, modeling sample-specific knowledge for downstream tasks through the integrated textual and visual structural knowledge.
As shown in Figure 4 (a), the conditional adapter is developed based on CoOp [30]. For a given input image $x$, we utilize the feature $z$ of the input image and the relationship between the image prototype $\mathbf{V}$/text prototype $\mathbf{W}$ to generate a sample-specific residual. This entails processing both the feature $z$ and the prototypes through a transformer block, which then produces the sample-specific residual. Specifically, we apply the image feature as query and the prototype embeddings as both key and value into an attention block. Then, the visual residual $R_{v}\in\mathbb{R}^{d\times C}$ is calculated as an exemplar of the sample-specific residual using the following set of equations:
@@ 4.1 Conditional Adapter
$R_{v}={\rm Norm}({\rm FFN_{0}}({\rm Softmax}(\frac{Q_{v}\cdot K_{v}^{\top}}{\sqrt{d}})V_{v})+{\rm FFN_{1}}(x)),$ (4)
@@
4.2 Nearest Neighbor Rectification
Inspired by C2A [21] and SuS-X [23] which enhance prototypes using generative diffusion models [20, 3, 4] or expansive external databases, our work seeks to enrich the limited knowledge obtained from few-shot samples. We aim to leverage the potential of unlabeled samples, circumventing the need for auxiliary or synthetic data, to mitigate the inherent biases and constraints associated with few-shot learning scenarios.
@@ 4.2 Nearest Neighbor Rectification
4.3 Tuning for Downstream Tasks
To ensure that the fused features $P$ we obtain do not significantly deviate from those generated by the pre-trained CLIP model, and to enhance the generalizability of model, we employ a consistency constraint. This constraint encourages the fused features to remain similar to the representations from the original pre-trained model, which has learned robust and generalizable features from a vast amount of data. We use a pre-trained large language model (LLM) to generate more descriptive sentences from template textual inputs, which are then used to guide the fused features. By doing so, we ensure that the fused features have a high degree of similarity with the textually-augmented representations provided by the LLM. We can denote the consistency constraint as:

### v2

@@ Introduction
Leveraging the vast array of image-text association pairs, the trained visual-language models (VLMs) embody critical general knowledge, thereby exhibiting enhanced generalizability across diverse tasks. In recent developments, several visual-language models like Contrastive Language-Image Pre-training (Radford et al. 2021) (CLIP), Flamingo (Alayrac et al. 2022), and ALIGN (Jia et al. 2021) have been introduced. While VLMs are proficient at extracting visual and textual descriptions, their training necessitates extensive, high-quality datasets. Nevertheless, amassing substantial data for training task-specific models in real-world visual-language applications poses a challenge. To circumvent this issue, efficient transfer learning (ETL) methods (Zhou et al. 2022b; Yao, Zhang, and Xu 2023; Zhou et al. 2022a; Li et al. 2023; Zhu et al. 2023b)have been devised to fine-tune pre-trained VLMs for downstream tasks, demonstrating remarkable efficacy in various few-shot or zero-shot visual recognition tasks.
@@
There are two notable ETL methods for VLMs: prompt tuning (Zhou et al. 2022b; Yao, Zhang, and Xu 2023; Zhou et al. 2022a; Zhu et al. 2023a) and adapter-style tuning (Gao et al. 2023; Li et al. 2023; Zhu et al. 2023b). Prompt tuning modifies the textual classifier by incorporating learnable prompts at the input, significantly outperforming zero-shot CLIP with few-shot examples, seen in CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a). Conversely, adapter-style tuning adjusts textual classifiers or visual features at the output, using the textual encoder just once. For example, CLIP-Adapter (Gao et al. 2023) employs a bottleneck layer to fine-tune embeddings, surpassing zero-shot CLIP by 3.02% on ImageNet in a one-shot scenario. TaskRes (Yu et al. 2023) adjusts textual embeddings with task-specific parameters, acting as independent residuals. Another approach enhances downstream task knowledge using large scale databases or generative models, like Sus-X (Udandarao, Gupta, and Albanie 2023) and C2A (Roy et al. 2023).
@@
However, two main limitations exist in current ETL research: 1) Many approaches, like CLIP-Adapter (Gao et al. 2023) and TaskRes (Yu et al. 2023), focus on task-specific knowledge from only one modality, relying solely on visual or textual features for adaptation. 2) Using generative models or large scale datasets to augment datasets typically requires substantial computational resources, impacting efficiency. In low-data scenarios, limited samples struggle to adequately reveal structural knowledge for downstream tasks, potentially biasing the model toward certain attributes like color or shape. This can limit transferability and generalization. Therefore, integrating multi-modality knowledge and discovering efficient strategies to enrich the dataset are essential for VLM tuning, addressing the biases that stem from data scarcity.
@@
To alleviate these limitations, we propose two prototype rectification strategies, referred to as conditional prototype rectification learning, which aim to model task-specific knowledge for downstream tasks by integrating textual and visual structural knowledge as well as knowledge from unlabeled samples. To tackle the first limitation, we introduce the Conditional Adapter (CoAdapter), which leverages the input image alongside visual and textual prototypes to capture structured knowledge pertinent to downstream tasks, producing sample-specific prototypes. This equips the feature adapter with a comprehensive and accurate understanding essential for these tasks. Addressing the second limitation, we devise the Nearest Neighbor Rectification (NNR) strategy. It identifies the $k$ nearest samples from the unlabeled dataset that align with the prototypes generated by the CoAdapter, facilitating rectification. This process effectively utilizes the latent knowledge within the unlabeled samples, enriching the model’s learning context. As shown in Figure 1, CPR achieves significant improvements over the previous state-of-the-art algorithms on generalization tasks.
@@
  • We introduce a novel adapter-style tuning approach, the Conditional Adapter (CoAdapter), for efficiently transfer learning of vision-language models. This strategy harnesses dual-modality structure knowledge from both textual and visual domains, allowing the feature adapter to utilize integrated visual and textual insights. This fusion facilitates enhanced learning of task-specific knowledge from downstream tasks, leading to effective VLM tuning.
@@
  • We evaluate our CPR across 11 popular benchmarks in both few-shot classification and base-to-new generalization settings. The evaluations reveal that CPR markedly surpasses earlier prompt-based and adapter-style methodologies, showcasing notable efficacy even in demanding fine-grained image classification tasks, exemplified by Caltetch101.
@@
Related Works
Visual-Language Pre-training
@@
The principal aim of Visual-Language Pre-training is to execute multimodal pre-training employing extensive datasets comprising image-text pairs, aiming to develop universal representations for text and images. This approach is intended to improve the effectiveness of model in various downstream tasks. These tasks notably include zero-shot learning, few-shot classification, cross-modal generation, open-world segmentation, and open-world detection, among others. In this context, models such as VisualBERT (Li et al. 2019), OSCAR (Li et al. 2020), and Uniter (Chen et al. 2020) use BERT (Devlin et al. 2019) to process raw text, achieving significant results in multimodal tasks like Visual Question Answering (VQA), which demonstrates their substantial application potential. These successes emphasize the importance of well-designed fusion encoders in merging cross-modal interactions. Additionally, recent studies like CLIP (Radford et al. 2021), DeCLIP (Li et al. 2022), and ALIGN (Jia et al. 2021) have demonstrated the effectiveness of visual-language contrastive learning in producing features that are useful for downstream tasks. These features are acquired by simply calculating the dot product between visual and linguistic embeddings, offering a straightforward understanding of multimodal interactions. Moreover, these approaches avoid the need for self-attention or cross-attention mechanisms, allowing for the precomputation and storage of multimodal embeddings without the need for further modifications, thus increasing efficiency and easing their integration into other tasks.
Efficient Transfer Learning
@@
Efficient Transfer Learning aims to transfer task knowledge to downstream tasks by adjusting parameters in pre-trained visual-language models. ETL methods include prompt tuning and adapter-style tuning.
Prompt tuning uses task-related textual tokens to infer task-specific knowledge, exemplified by the text template of CLIP for zero-shot prediction. Context Optimization (CoOp) (Zhou et al. 2022b) uses learnable soft prompts from few-shot examples, but generates fixed prompts for each task, ignoring image feature variation. Conditional Context Optimization (CoCoOp) (Zhou et al. 2022a) creates image-specific prompts using a lightweight neural network. ProDA (Derakhshani et al. 2023) learns a prior distribution for prompts, and ProGrad (Zhu et al. 2023a) selectively updates prompt gradients aligned with ”general knowledge.”
Adapter-style tuning modulates textual and visual features for task adaptation. CLIP-Adapter (Gao et al. 2023) uses a bottleneck layer for few-shot classification, and TaskRes (Yu et al. 2023) introduces a task-independent adapter.
@@
In existing methods, the works most relevant to our study are CoCoOp and Sus-X (Udandarao, Gupta, and Albanie 2023). CoCoOp can be regarded as the baseline model for our proposed CPR. Unlike CoCoOp, our CPR considers not only the input image itself but also the structural relationship between the input image and both visual and textual prototypes to further enhance the performance of learnable prompts. Sus-X introduces a text-image generative model to augment the original support set, whereas CPR utilizes unlabeled samples from the test set for prototype refinement, offering greater efficiency.
Preliminaries
@@
We first provide brief reviews on the adopted VLM, i.e., contrastive language-image pre-training (CLIP) (Radford et al. 2021), recap two mainstream approaches for ETL on VLMs, i.e., prompt tuning and adapter-style tuning.
Contrastive Language-Image Pre-training
@@
Contrastive Language-Image Pre-training (Radford et al. 2021) aims to acquire visual representations through natural language supervision. CLIP consists of two encoders: one for images and another for text. The image encoder can be either ResNet (He et al. 2016) or Vision Transformer (Dosovitskiy et al. 2021) (ViT), transforming images into feature vectors. The text encoder is a Transformer (Vaswani et al. 2017) that takes a sequence of word tokens as input and produces a vectorized representation.
@@
CLIP is trained using 400 million image-text pairs and employs a contrastive loss to learn a joint embedding space for both modalities, enabling CLIP to effectively capture a wide range of visual concepts and learn universal visual representations. Specifically, for an input image $x$ belonging to one of the classes $Y=\{y_{1},y_{2},...,y_{C}\}$, the image encoder $f$ extracts image features $z=f(x)\in\mathbb{R}^{d}$. To obtain the corresponding text features $w_{i}\in\{1,...,C\}$, names of all given classes are filled into the fixed prompt template ”a photo of a [CLASS]” (e.g., bear, urocissa erythrorhyncha), generating text descriptions $A$, which are then further encoded into text embeddings $W=g(A)\in\mathbb{R}^{d\times C}$. Image-text alignment is optimized based on the cosine similarity of features.
@@
Revisiting Previous Efficient Transfer Paradigms
@@
CoOp (Zhou et al. 2022b) introduces prompt tuning to VLMs for the first time, utilizing $M$ learnable context vectors $v=\{v_{1},v_{2},...,v_{M}\}$to generate task-specific templates, replacing the fixed text prompt templates used in CLIP. Specifically, $t_{i}=\{v_{1},v_{2},...,v_{M},c_{i}\}$ integrates learnable context vectors $v$ with the class token embedding $c_{i}$. This combined prompt is then input into the text encoder $g(\cdot)$. During training process, CoOp maintains the pre-trained VLMs parameters unchanged and solely adjusts the learnable vectors $v$. CoCoOp (Zhou et al. 2022a) is an enhanced version of CoOp. CoOp suggests that contexts conditioned on individual instances offer superior generalization because they pivot attention from a narrow set of classes (thus reducing overfitting risks) toward each unique input instance, facilitating broader task applicability. Meanwhile, CoCoOp introduces a Meta-Net to produce a conditional token that is subsequently amalgamated with the context vectors, enhancing adaptability and specificity to each instance.
@@
where $\alpha$ is a scaling factor.In CLIP-Adapter, the adapter module $\phi_{\omega}$ consists of two linear transformation layers with a ReLU activation function situated between them. CLIP-Adapter explores the use of adapters for both visual and textual components, applying the adapter module separately to the image and text branches of CLIP. This investigation reveals that both types of adapters—visual and textual—yield similarly effective performance. When training on downstream tasks, adapter-style approaches like CLIP-Adapter focus on tuning the parameters of these adapter modules only, leaving the foundational pre-trained CLIP model parameters unchanged. This strategy allows for task-specific customization while maintaining the general capabilities of the underlying model. Moreover, recent developments in SuS-X (Udandarao, Gupta, and Albanie 2023) and C2A (Roy et al. 2023) have utilized generative diffusion models or large-scale external databases to achieve enhanced feature representations $f^{\prime}$.
Approach
@@
Conditional Adapter
Previous works like CoCoOp (Zhou et al. 2022a) considered how input images impact text prototypes. And CLIP-Adapter (Gao et al. 2023), TaskRes (Yu et al. 2023), and Tip-Adapter (Zhang et al. 2022) derive features using separate visual or textual features. In contrast, we propose conditional adapter (CoAdapter), which leverages the connections between input images and both visual and textual prototypes, modeling sample-specific knowledge for downstream tasks through the integrated textual and visual structural knowledge.
As shown in Figure 4 (a), the conditional adapter is developed based on CoOp (Zhou et al. 2022b). For a given input image $x$, we utilize the feature $z$ of the input image and the relationship between the image prototype $\mathbf{V}$/text prototype $\mathbf{W}$ to generate a sample-specific residual. This entails processing both the feature $z$ and the prototypes through a transformer block, which then produces the sample-specific residual. Specifically, we apply the image feature as query and the prototype embeddings as both key and value into an attention block. Then, the visual residual $R_{v}\in\mathbb{R}^{d\times C}$ is calculated as an exemplar of the sample-specific residual using the following set of equations:
@@
$R_{v}={\rm Norm}({\rm FFN_{0}}({\rm Softmax}(\frac{Q_{v}\cdot K_{v}^{\top}}{\sqrt{d}})V_{v})+{\rm FFN_{1}}(z)),$ (4)
@@
Nearest Neighbor Rectification
Inspired by C2A (Roy et al. 2023) and SuS-X (Udandarao, Gupta, and Albanie 2023) which enhance prototypes using generative diffusion models (Rombach et al. 2022; Brooks, Holynski, and Efros 2023; Chen et al. 2023) or expansive external databases, our work seeks to enrich the limited knowledge obtained from few-shot samples. We aim to leverage the potential of unlabeled samples, circumventing the need for auxiliary or synthetic data, to mitigate the inherent biases and constraints associated with few-shot learning scenarios.
@@
Tuning for Downstream Tasks
To ensure that the fused features $P$ we obtain do not significantly deviate from those generated by the pre-trained CLIP model, and to enhance the model’s generalization ability, we employ a consistency constraint. This constraint encourages the fused features to remain similar to the representations from the original pre-trained model, which has learned robust and generalizable features from a vast amount of data. We use a pre-trained large language model (LLM) to generate more descriptive sentences from template textual inputs, which are then used to guide the fused features. By doing so, we ensure that the fused features have a high degree of similarity with the textually-augmented representations provided by the LLM. We can denote the consistency constraint as:
@@
Dataset CLIP CoOp CoCoOp KgCoOp MaPLe PromptSRC TCP CPR
(ICML’21) (IJCV’22) (CVPR’22) (CVPR’23) (CVPR’23) (ICCV’23) (CVPR’24)
Base 69.34 82.63 80.47 80.73 82.28 84.12 84.13 84.02
Average New 74.22 67.99 71.69 73.60 75.14 75.02 75.36 76.22
HM 71.70 74.60 75.83 77.00 78.55 79.31 79.51 79.93
Base 72.43 76.46 75.98 75.83 76.66 77.75 77.27 77.31
ImageNet New 68.14 66.31 70.43 69.96 70.54 70.70 69.87 70.72
HM 70.22 71.02 73.10 72.78 73.47 74.06 73.38 73.87
Base 96.84 98.11 97.96 97.72 97.74 98.13 98.23 98.64
Caltech101 New 94.00 93.52 93.81 94.39 94.36 93.90 94.67 95.43
HM 95.40 95.76 95.84 96.03 96.02 95.97 96.62 97.01
Base 91.17 94.24 95.20 94.65 95.43 95.50 94.67 95.58
OxfordPets New 97.26 96.66 97.69 97.76 97.76 97.40 97.20 97.79
HM 94.12 95.43 96.43 96.18 96.58 96.44 95.92 96.67
Base 63.37 76.20 70.49 71.76 72.94 78.40 80.80 78.32
StanfordCars New 74.89 69.14 73.59 75.04 74.00 74.73 74.13 75.12
HM 68.65 72.49 72.01 73.36 73.47 75.52 77.32 76.69
Base 72.08 97.63 94.87 95.00 95.92 97.90 97.73 97.15
Flowers102 New 77.80 69.55 71.75 74.73 72.46 76.77 75.57 79.07
HM 74.83 81.23 81.71 83.65 82.56 86.06 85.23 87.18
Base 90.10 89.44 90.70 90.50 90.71 90.63 90.57 90.95
Food101 New 91.22 87.50 91.29 91.70 92.05 91.50 91.37 92.09
HM 90.66 88.46 90.99 91.09 91.38 91.06 90.97 91.52
Base 27.19 39.24 33.41 36.21 37.44 42.30 41.97 41.26
FGVCAircraft New 36.29 30.49 23.71 33.55 35.61 36.97 34.43 35.78
HM 31.09 34.30 27.74 34.83 36.50 39.46 37.83 38.33
Base 69.36 80.85 79.74 80.29 80.82 82.83 82.63 82.77
SUN397 New 75.35 68.34 76.86 76.53 78.70 79.00 78.20 79.48
HM 72.23 74.07 78.27 78.36 79.75 80.87 80.35 81.09
Base 53.24 80.17 77.01 77.55 80.36 82.60 82.77 82.95
DTD New 59.90 47.54 56.00 54.99 59.18 57.50 58.07 58.57
HM 56.37 59.68 64.85 64.35 68.16 67.80 68.25 68.66
Base 56.48 91.54 87.49 85.64 94.07 92.40 91.63 92.93
EuroSAT New 64.05 54.44 60.04 64.34 73.23 68.43 74.73 74.57
HM 60.03 68.27 71.21 73.48 82.35 78.63 82.32 82.68
Base 70.53 85.14 82.33 82.89 83.00 86.93 87.13 86.35
UCF101 New 77.50 64.47 73.45 76.67 78.66 78.33 80.77 79.83
HM 73.85 73.37 77.64 79.65 80.77 82.41 83.83 82.96